OpenStudy (anonymous):

Determine whether Rolle's theorm can be applied on the closed interval [a, b]. If Rolle's theorm can be applied, find all vales of c in the interval [a,b] such that f'(c)=0. If Rolle's theorm cannot be applied, explain why not.

OpenStudy (anonymous):

f(x)=(x-1)(x-2)(x-3), [1,3]

OpenStudy (anonymous):

since \[f(1)=f(3)=0\] and since any polynomial is differentiable you can use rolle

OpenStudy (anonymous):

\[f(x)=x^3-6 x^2+11 x-6\] \[f'(x)=3x^2-12x+11\] so rolle says there somewhere in [1,3] there is a number c with \[f'(c)=0\] set \[3x^2-12x+11=0\] solve for x

OpenStudy (anonymous):

that is where i am stuck, i got that far

OpenStudy (anonymous):

use the quadratic formula if it does not factor

OpenStudy (anonymous):

if that is too annoying, then cheat http://www.wolframalpha.com/input/?i=3x^2-12x%2B11%3D0

OpenStudy (anonymous):

-b+-sqrt(b^2-4ac)/2a?

OpenStudy (anonymous):

yup